OpenStudy (anonymous):

If Benito earns $187.50 for 12.5 hours of work, how many hours must he work to earn $465.00

OpenStudy (anonymous):

all you gotta do is use a proportion.... do you know what thoes are and how they work?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

well a proportion is just two fractions that a;ways equal eachother... for example \[\frac{ 1 }{ 2 }=\frac{ 1 }{ x }\] all you have to do is cross multiply and solve for x so for this one \[1*x=2*1 and x=2\]

OpenStudy (anonymous):

so for you you jest need to make a proportion with your data... it should looke something like this \[\frac{ 187.5 }{ 12.5 }=\frac{ 465 }{ x }\] now you always have to make sure that the two peices of data that go together are in the same part of the fraction. notice how i have both dollar amounts in the numerator and both hour amounts in the denominator

OpenStudy (anonymous):

do you understand so far?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so try cross multiplying it and tell me what you get

OpenStudy (anonymous):

How do I suppose to cross multiply 187.5 and x

OpenStudy (anonymous):

you just get 187.5x and that equals 465*12.5

OpenStudy (anonymous):

then to solve for x you need to divide 465*12.5 by 187.5 and you should end up with your answer :)

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so what did you get?